# Future-Proofing Class Incremental Learning

### v1

@@ 4.3.3 Other classifiers
Similarly to FeTrIL, the performance of NMC is highly dependent on the quality of the features produced by the backbone. It struggles with the Learning From Scratch procedure where the number of classes and data samples used to train the feature extractor is limited. Our method improves the performance of the NMC in every considered setting and achieves the most significant improvements with the Learning From Scratch procedure by reaching up to three times the final accuracy of the baseline method on CIFAR100 B0 Inc5.
4.3.4 Quality and diversity
@@ 5 Conclusion
In this work, we introduced a new method for Exemplar-Free Class Incremental Learning which leverages large pre-trained diffusion model to generate images from future classes. Experimental results highlight how our method can significantly improve the accuracy of state-of-the-art approaches while only modifying the initial step. We found that our method requires less data than traditional methods relying on real curated datasets. While we limit our present study to feature extractors trained from scratch during the first incremental step, in future work, we will further investigate how synthetic images of future classes can be used to adapt a general pre-trained foundation model. Moreover, we will also explore how synthetic data of future classes can be leveraged during the whole training procedure, not only the first incremental step.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment

### v2

@@ 4.3.6 Comparison with additional baselines
Additionally to FeTrIL [7], we report in Table 11 the performance of our method combined with a Nearest Class Mean (NCM) Classifier and FeCAM [33]. NCM is a simple yet effective method [34, 29] for Class-Incremental Learning with a fixed feature extractor that does not require memory. FeCAM is a recently proposed approach for Exemplar-Free Class-Incremental Learning achieving new state-of-the-art performances. Similarly to FeTrIL, the performance of both NCM and FeCAM is highly dependent on the quality of the features produced by the backbone. It struggles with the Learning From Scratch procedure where the number of classes and data samples used to train the feature extractor is limited. Our method improves the performance of the NCM classifier and FeCAM in every considered setting and achieves the most significant improvements with the Learning From Scratch procedure on CIFAR100 B0 Inc5, reaching a final accuracy up to two times higher than the baseline approach.
Furthermore, in Table 12 we compare our approach with the concurrently released method DiffClass [71] which leverages the Stable Diffusion model for generating samples from past classes. While our method takes advantage of a larger backbone and can achieve higher performances as detailed in the ablation study, the comparison on CIFAR100 is done using ResNet-32 for the backbone as the results for DiffClass using ResNet-18 are not available. Experimentally, DiffClass achieves higher performance on CIFAR100 but is outperformed by our approach on ImageNet-Subset. It should be noted that Meng et al. [71] proposed several new components to mitigate the distribution gap between real and synthetic samples [47]. Some of those contributions, such as a multi-distribution matching technique to finetune the diffusion model and a selective synthetic image augmentation technique, could be adapted and combined with our proposed approach FPCIL to further improve its performance. Finally, our proposed method FPCIL conducts the computationally expensive task of generating synthetic samples using diffusion model only during the first incremental step, before freezing the feature extractor and only updating the classifier afterward. This allows us to potentially conduct the first incremental step on specialised hardware in a data center before deploying the model on devices with limited hardware where each following incremental step could be conducted in a matter of minutes on a CPU using approaches such as FeTrIL or FeCAM. In comparison, DiffClass finetunes the diffusion model, generates synthetic samples of past classes, and finetunes the whole feature extractor at each incremental step, resulting in much larger computational cost and requiring specialised hardware during whole training procedure. This makes our approach preferable if applicable.
@@ 5 Conclusion
In this work, we introduced a new method for Exemplar-Free Class-Incremental Learning which leverages large pre-trained diffusion model to generate images from future classes. Experimental results highlight how our method can significantly improve the accuracy of state-of-the-art approaches while only modifying the initial step. We found that our method requires less additional data than traditional methods relying on real curated datasets. While we focused our present study to feature extractors trained from scratch during the first incremental step, in future work, we will further investigate how synthetic images of future classes can be used to adapt a general pre-trained foundation model. Moreover, we will also explore how synthetic data of future classes can be leveraged during the whole training procedure, not only during the first incremental step.
[…]
Acknowledgements
@@
[…]
Credit Author Statement Quentin Jodelet: Conceptualization, Investigation, Software, Writing - Original Draft. Xin Liu: Conceptualization, Supervision, Resources, Funding acquisition, Writing - Review & Editing. Yin Jun Phua: Conceptualization, Supervision, Writing - Review & Editing. Tsuyoshi Murata: Supervision, Resources, Funding acquisition, Writing - Review & Editing.
[…]
Data Availability
Benchmark datasets CIFAR-100, and ImageNet are publicly available. Diffusion models used for generating synthetic datasets are either publicly available or behind a paid API.
Declarations
[…]
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.